smithkids:

A newspaper reports the following changes in the temperature of a city over 4 days: 1.34 degrees Celsius, negative 5 over 7 degrees Celsius, −0.75 degrees Celsius, 4 over 9 degrees Celsius Write an expression to show the average daily change in the temperature of the city. Write the steps to solve the expression. Round your answer to the nearest hundredth.

jaylaa1:

so first u wanna see that the total of change in the temperature of the city Change of temp = 1.34 - 5/7 - 0.75 + 4/9 And now u would do = 1.34 - 0.75 + (4/9 - 5/7) Next,=0.61 + (28/63 - 45/63),=0.61 - 17/63 Last = 21.43/63 (21.43/63)÷4= 0.85